# MoGe: Unlocking Accurate Monocular Geometry Estimation for Open-Domain Images with Optimal Training Supervision

### v2

@@ Appendix C More Results
LeReS 7.33 95.4 12.4 86.7 10.2 90.1 8.45 92.9 4.33 99.7 28.9 59.6 23.3 73.1 10.8 88.3 13.2 85.7 8.25
ZoeDepth 5.24 97.7 5.86 95.6 8.08 94.0 6.22 96.1 2.59 99.9 26.9 66.3 14.0 81.8 8.17 92.0 9.63 90.4 6.75
DUSt3R 4.27 98.1 7.67 92.2 5.60 96.2 4.47 96.7 2.64 99.8 39.9 56.8 17.4 76.3 7.09 92.8 11.1 88.6 6.75
Metric3D V2 13.4 81.4 3.69 98.3 4.38 97.6 8.56 92.3 1.81 100 21.8 72.5 7.32 94.2 7.62 90.4 8.57 90.8 5.29
UniDepth 3.80 98.7 3.64 98.7 5.34 97.2 4.04 98.1 2.55 99.9 28.5 60.9 9.83 89.3 5.94 95.5 7.96 92.3 3.62
MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 5.00
DA V1 4.24 98.4 5.40 97.0 4.70 98.2 4.20 97.6 1.54 100 20.4 77.5 12.6 87.0 5.72 95.7 7.36 93.9 3.00
DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 3.06
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
Table 6: Full table of comparison for depth map estimation. *: methods that have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results for each version (smaller font size). Gray numbers denote models trained on respective benchmarks.
More qualitative comparisons.
@@ Appendix D Limitations and Future Work
Figure 9: A failure case. Our model does not capture the thin structures of the fence, leading to a flattened geometry.

### v3

@@ Appendix C More Results
LeReS 7.31 95.5 12.2 87.1 10.2 90.1 8.44 92.9 4.33 99.7 28.9 59.6 23.4 73.0 10.7 88.3 13.2 85.8 8.25
ZoeDepth 5.21 97.7 5.84 95.6 8.07 94.0 6.19 96.1 2.60 99.9 26.9 66.3 14.1 81.7 8.17 92.0 9.63 90.4 6.75
DUSt3R 4.24 98.1 7.72 92.1 5.60 96.2 4.49 96.6 2.63 99.8 40.0 56.7 17.4 76.2 7.10 92.8 11.1 88.6 6.75
Metric3D V2 13.4 81.5 3.76 98.2 4.30 97.7 8.55 92.3 1.80 100 21.8 72.4 7.35 94.1 7.70 90.2 8.58 90.8 5.29
UniDepth V1 3.78 98.7 3.64 98.7 5.34 97.2 4.06 98.1 2.56 99.9 28.6 60.7 9.94 89.1 5.95 95.5 7.98 92.2 3.62
MiDaS V3.1 4.58 98.1 6.25 94.7 5.77 96.8 4.73 97.4 1.86 100 21.3 73.1 14.5 82.6 6.05 94.9 8.13 92.2 5.00
DA V1 4.20 98.4 5.40 97.0 4.68 98.2 4.18 97.6 1.54 100 20.1 77.6 12.7 86.9 5.69 95.7 7.31 93.9 3.00
DA V2 4.14 98.3 5.61 96.7 4.71 97.9 3.47 98.5 1.24 100 21.4 72.8 13.1 86.4 5.29 96.1 7.37 93.3 3.12
Ours 3.38 98.6 4.05 98.1 3.11 98.9 3.23 98.0 0.96 100 18.4 79.5 8.99 91.5 3.98 97.2 5.76 95.2 1.44
Table 6: Full table of comparison for depth map estimation. * methods have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results in smaller text size for each version. Gray numbers denote models trained on respective benchmarks.
More qualitative comparisons
@@ Appendix D Limitations and Future Work
Figure 9: A failure case. Our model fails to capture the thin structure of the fence, leading to a flattened geometry.